OpenStudy (anonymous):

Use the distributive property to rewrite the expression and solve. 6 94 = 6 ? + 6 4 = ?

OpenStudy (anonymous):

I will give MEDAL AND FAN

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@Destinymasha

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@periy

OpenStudy (anonymous):

lol no one is coming

OpenStudy (anonymous):

@yrelhan4

OpenStudy (anonymous):

@samson245

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

Wait I messed up thequestion

OpenStudy (anonymous):

Use the distributive property to rewrite the expression and solve. 6 * 94 = 6 * ? + 6 * 4 = ?

OpenStudy (mathmate):

Hints: Distributive property: a(b+c)=ab+ac Substitute a,b and c with appropriate numbers.

OpenStudy (anonymous):

There

OpenStudy (anonymous):

I dont see what you mean

OpenStudy (anonymous):

I dont get it......

OpenStudy (perl):

6 * 94 = 6 ( ? + 4) = 6 * ? + 6 * 4